graph the line which passes through the point (6,-1) and is perpendicular to the line with equation 2x-6y=-14
Answer by stanbon(75887)

graph the line which passes through the point (6,-1) and is perpendicular to the line with equation 2x-6y=-14
---
The given line has a slope = 1/3
---
Any perpendicular line must have a slope of -3.
----------------
Form: y = mx + b
You know m = 1/3 and y = -1 when x = 6; solve for "b":
-1 = (1/3)(6) + b
-1 = 2 + b
b = -3
-------------------
Ans: y = -3x -3
=======================
Cheers,
Stan H.